OpenStudy (anonymous):

An original data set has the values eleven, 8, 7, twelve, and 7. What would need to be added to each value to change the mean to fifteen?

OpenStudy (mathmate):

Hint: Shifting all data by x will shift the mean by the same amount. The value needed to shift would be the difference between 15 and the mean of the current data.

OpenStudy (anonymous):

15-9=6, So the answer would be 6? is that right @mathmate

OpenStudy (mathmate):

The method is correct if 9 is the mean of the current dataset. Does the mean of the current dataset equal nine?

OpenStudy (anonymous):

11+8+7+12+7=45 45/5= 9 15-9= 6 So the answer is 6?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Sorry, I didn't see the eleven, but I saw the seven. Yes, you got it right! :)

OpenStudy (mathmate):

*twelve

OpenStudy (anonymous):

Also, Randall set a personal goal to exercise at least forty-five minutes for half of the days during the last two months. He recorded his exercise time each day. Use the box and whisker plot of his data to determine whether he reached his goal. Explain.

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

 The box-and-whisker plot shows clearly that half of the days he worked out less than 30 minutes, and half of the days he worked out more than 30 minutes (= median). Looking at the Q3 number (45 minutes), that means 3/4 of the days he worked out less than 45 minutes, and 1/4 of the days the time is above 45 minutes. You'll make your conclusions accordingly.